OpenStudy (emilyjones284):

ill give a medal (: If f(x)=x+3 over 4, what is the equation for f^–1(x)?

OpenStudy (anonymous):

think of it as resolving the equation again i.e we have \[y = \frac{x + 3}{4}\] if we switch the y and the x we can solve for y again so now we have \[x = \frac{y + 3}{4}\] now all you have to do is solve for y again and you will have your answer :)

OpenStudy (emilyjones284):

I dont understand the solving part haha

OpenStudy (anonymous):

ah i see we have\[x = \frac{y+3}{4}\] since we cant do anything until the fraction part is over we have to muliply by 4 on both side which will give us \[4x =y +3\] now we only have one more step until we can finsh which is adding 3 from both sides now we have \[4x + 3 = y\]or in our case\[f^{-1}(x)=4x +3\]

OpenStudy (emilyjones284):

Ok thanks a lot I got it now (:

OpenStudy (anonymous):

no problem :P